The James Company is expected to pay a total of $1.86 in dividends next year. Their dividends are expected to increase by 3% annually. James Company stock is selling for $17.54 a share. What is the capitalization rate on this stock?
Answer #1

Price = D1/(ke-g) = $17.54 D1 = $1.86 3.00% ke - g = 10.60% (1.86/17.54] Capitalization rate = ke = 13.60% 13% +10.60% ]